there are 17 balls in a basket. The number of red balls is 3 more than the number of blue balls. The number of yellow balls is 4 less than the number of red balls. How many blue balls are in the basket? how many red balls? How many yellow balls?

Answer by Alan3354(69443)

there are 17 balls in a basket. The number of red balls is 3 more than the number of blue balls. The number of yellow balls is 4 less than the number of red balls. How many blue balls are in the basket? how many red balls? How many yellow balls?
--------
r = b + 3
y = r - 4 --> y = b - 1
r + b + y = 17
-----
Sub for r & y in the 3rd equation
---
b+3 + b + (b-1) = 17
3b + 2 = 17
3b = 15
b = 5
============
r = 8
============
y = 4